Circular No. 7278

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

XTE J1859+226 G. G. Pooley, Mullard Radio Astronomy Observatory; and R. M.

Hjellming, National Radio Astronomy Observatory, report: “We have made Ryle Telescope (RT) and Very Large Array (VLA) observations of a radio counterpart to the x-ray transient XTE J1859+226 (IAUC 7274). On Oct. 11.88 UT, the RT detected a roughly 10-mJy radio source at 15 GHz with α = 18^{h}58^{m}42^{s}.0, δ = +22^{o}39′10′′ (equinox 2000; uncertainties 0^{s}.5 in α and 30′′ in δ). A further RT observation on Oct. 12.86 gave 12.5 mJy (uncertainty 1.5 mJy). The VLA observations on Oct. 13.06 confirm the presence of a radio source with preliminary flux densities of 8.7, 10.8, 11.1, 11.2, and 11.3 mJy at frequencies of 1.4, 4.9, 8.5, 14.9, and 22.5 GHz, with position end figures of 41^{s}.49, 29′′.8 (uncertainty 0′′.3). This is consistent with the position of the optical counterpart reported on IAUC 7276. No radio source was present at this position in the NVSS survey in Feb. 1995 with an upper limit of ∼ 3 mJy.”

COMET C/1999 H1 (LEE)

Ephemeris extension to IAUC 7266:

1999TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Oct. 19 23^{h}02.^{m}06 +28^{o}50.^{′}9 1.036 1.908 139.^{o}8 19.^{o}7  9.4
24 22 48.44 +23 22.9 1.147 1.978 134.8 20.9  9.8
29 22 39.04 +18 51.4 1.272 2.047 128.9 22.2 10.1
Nov. 3 22 32.62 +15 09.8 1.408 2.115 122.7 23.3 10.5
8 22 28.38 +12 09.5 1.552 2.183 116.6 24.0 10.8
13 22 25.77 + 9 42.9 1.702 2.251 110.7 24.3 11.2
18 22 24.39 + 7 43.5 1.856 2.317 105.0 24.3 11.5
23 22 23.97 + 6 06.0 2.012 2.384  99.5 24.1 11.8
28 22 24.30 + 4 46.4 2.170 2.449  94.2 23.7 12.1
Dec. 3 22 25.23 + 3 41.2 2.329 2.515  89.1 23.1 12.3
8 22 26.66 + 2 48.1 2.486 2.579  84.1 22.3 12.6

Visual m_{1} estimates: Sept. 12.17 UT, 7.7 (J. Bortle, Stormville, NY, 20×80 binoculars); 20.78, 7.9 (Y. Nagai, Yamanashi, Japan, 11×80 binoculars); Oct. 5.16, 7.8 (R. Keen, Mt. Thorodin, CO, 7×50 binoculars); 7.17, 8.4 (A. Pereira, Cabo da Roca, Portugal, 9×34 binoculars); 8.94, 9.2 (K. Sárneczky, Ráktanya, Hungary, 20×60 binoculars); 11.22, 8.5 (Keen, 0.15m refl.); 11.80, 8.6 (B. H. Granslo, Fjellhamar, Norway, 0.07-m refractor).

1999 October 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green